OpenStudy (hanamalik):

Find the distance between the points (2, 3) and (0, 6).

OpenStudy (fibonaccichick666):

so do you know the distance formula?

pooja195 (pooja195):

looks like this :-) \[\huge d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\]

OpenStudy (fibonaccichick666):

if not, we can derive it.

OpenStudy (hanamalik):

\[\sqrt{?} (x2-x1)^{2}+(y2-y1)^{2}\]

OpenStudy (fibonaccichick666):

yep yep, so now all you need to do is pick your \((x_1,y_1)\) and the sub twos and plug them in

OpenStudy (fibonaccichick666):

go ahead, give it a try

OpenStudy (hanamalik):

2.45 units?

OpenStudy (fibonaccichick666):

leave it exact for me? That seems a little small

OpenStudy (fibonaccichick666):

remember to square the result of your subtractions

OpenStudy (hanamalik):

still lost.

OpenStudy (anonymous):

Answer would be the square root of 13 Label your points. (2,3) (0,6) 2 is X1 0 is X2 3 is Y1 6 is Y2 Plug them in and you'll get the equation d=√(0-2)^2 + (6-3)^2 Simplify d=√(-2)^2 + (3)^2 Square d=√4+9 d=√13